# On sharpened singular Adams’ type inequalities and applications to $(p,\frac{n}{2})$-biharmonic equations

### v2

@@ 2. Preliminary results
In this section, we shall discuss the basic properties of Schwarz symmetrization and some useful lemmas. In this regards, we refer to [7, 10]. First, we introduce about the symmetrization of Lebesgue measurable set $S$ in $\mathbb{R}^{n}$. Let $S^{\ast}$ be the open ball centered at the origin of $\mathbb{R}^{n}$ such that $|S^{\ast}|=|S|$. Further, we denote $|S|=0$ implies $S^{\ast}$ is the empty set, $|S|=+\infty$ implies $S^{\ast}=\mathbb{R}^{n}$ and $0<|S|<+\infty$ implies $S^{\ast}=B_{R}(0)$ such that $|S^{\ast}|=|S|=\sigma_{n}R^{n}$, where $\sigma_{n}$ is the volume of the unit ball in $\mathbb{R}^{n}$ and $R=R(S)$ is the volume radius of $S$. It is well known that $S^{\ast}$ is called the symmetric decreasing rearrangement of $S$. Let $f\colon S\to\mathbb{R}$ be a measurable function vanishing at infinity, which means that
@@
Its distribution function $\mu_{f}\colon[0,+\infty)\to[0,+\infty]$ is given by $\mu_{f}(t)=|\{x\in S\colon~{}|f(x)|>t\}|$ for all $t\geq 0$. Note that $\mu_{f}$ is always nonnegative, decreasing and right continuous function on $[0,+\infty)$. Moreover, the unidimensional rearrangement of $f$ is a function $f^{\sharp}\colon[0,|S|]\to[0,+\infty]$, which is defined by
@@
where $m_{\mu_{f}}$ is the distribution function of $\mu_{f}$. Besides this, $f^{\sharp}$ is a nonnegative, decreasing and right continuous function on $[0,+\infty)$. The maximal function of $f^{\sharp}$ is denoted by the symbol $f^{\sharp\sharp}$, and given by
$f^{\sharp\sharp}(s):=\frac{1}{s}\int_{0}^{s}f^{\sharp}(t)\,\mathrm{d}t,~{}\forall~{}~{}s>0.$ (2.1)
In addition, $f^{\sharp\sharp}$ is nonnegative, decreasing and continuous on $(0,+\infty)$ and $f^{\sharp}\leq f^{\sharp\sharp}$. Now we denote the function $f^{\ast}\colon S^{\ast}\to[0,+\infty]$ as the Schwarz symmetrization of $f$, alternatively known as the spherically symmetric decreasing rearrangement of $f$, and defined by
@@
In particular, $f^{\ast}$ is a radially symmetric nonnegative nonincreasing function. Moreover, if $\Psi\colon\mathbb{R}\to\mathbb{R}$ is nondecreasing, then we have
$(\Psi(f))^{\ast}=\Psi(f^{\ast}).$ (2.2)
@@
In addition, we also have $\|f\|_{p}=\|f^{\ast}\|_{p}$, $\forall~{}p\in[1,+\infty)$ and the following Hardy-Littlewood inequality holds.
@@
[7, Corollary 2.16] Suppose that $f$ and $g$ are two nonnegative measurable functions on $\mathbb{R}^{n}$ with $\mu_{f}(t)<+\infty$ and $\mu_{g}(t)<+\infty$ for all $t>0$. Then
$\int_{\mathbb{R}^{n}}f(x)g(x)\,\mathrm{d}x\leq\int_{\mathbb{R}^{n}}f^{\ast}(x)g^{\ast}(x)\,\mathrm{d}x$ (2.3)
with the understanding that when the left side is infinite, so is the right side. Moreover, if $\int_{\mathbb{R}^{n}}f(x)g(x)\,\mathrm{d}x<+\infty$, then the equality in (2.3) holds if and only if |$|\{(x,y)\in\mathbb{R}^{2n}\colon~{}f(x)<(f(y)~{}\text{and}~{}g(x)>g(y)\}|=0.$
@@
[41, Proposition 2.2] Let $u\in C^{\infty}_{0}(\mathbb{R}^{n})$ and $-\Delta u=f\in C^{\infty}_{0}(\mathbb{R}^{n})$, then there exists a constant $D(n)$ depends upon only $n$ such that
@@
[11, Lemma A.VI] If $u\in L^{\wp}(\mathbb{R}^{n}),~{}\wp\geq 1$, is a radial nonincreasing function, then there holds
@@
The following corollary follows from [49, Lemma 2.1].
@@
The function $s\mapsto\operatorname{\exp}(s)-\sum_{j=0}^{j_{0}-1}\frac{s^{j}}{j!}$ is continuous, increasing and convex in $[0,+\infty)$, where $j_{0}$ is defined in (1.3). Moreover, for any $\alpha>0$, $\wp>1$ and $n\geq 4$, we have
@@
[25, Theorem 1.2 and Lemma 3.5] If $q\in[1,+\infty]$, $r\in[1,+\infty)$, $j,~{}k\in\mathbb{N}$, $j<k$ such that $\frac{1}{p}=\frac{j}{kr}+\frac{k-j}{kq}$, then there exists a constant $C>0$ independent of $u$ such that
@@
3. Proof of Theorem 1.3 and Theorem 1.4
In this section, our aim is to prove Theorem 1.3 and Theorem 1.4. To establish Theorem 1.3, we shall use a rearrangement-free approach introduced by Lam-Lu [33], whereas the proof of Theorem 1.4 relies on conventional Schwarz symmetrization techniques.
@@
$H(u)<1\quad\text{and}\quad|S|\leq 2^{\frac{4p^{\ast}}{n(n-2)}}C,$ (3.1)
where $C>0$ is a suitable constant depending only on $n$ and $p$, thanks to the chain of embeddings $E\hookrightarrow D^{2,p}(\mathbb{R}^{n})\hookrightarrow L^{p^{\ast}}(\mathbb{R}^{n})$. This shows that $S$ is a bounded domain. In addition, we can notice that
@@
$J_{2}$$\leq\int_{\{x\in\mathbb{R}^{n}\colon~{}u(x)<1\}}\frac{1}{|x|^{\gamma}}\sum_{j=j_{0}}^{\infty}\frac{\alpha^{j}}{j!}|u|^{\frac{nj}{n-2}}\,\mathrm{d}x\leq\int_{\{x\in\mathbb{R}^{n}\colon~{}u(x)<1\}}\frac{1}{|x|^{\gamma}}\sum_{j=j_{0}}^{\infty}\frac{\alpha^{j}}{j!}|u|^{p^{\ast}}\,\mathrm{d}x$
$\leq\operatorname{\exp}(\alpha)\int_{\{x\in\mathbb{R}^{n}\colon~{}u(x)<1,~{}|x|<1\}}\frac{|u|^{p^{\ast}}}{|x|^{\gamma}}\,\mathrm{d}x+\operatorname{\exp}(\alpha)\int_{\{x\in\mathbb{R}^{n}\colon~{}u(x)<1,~{}|x|\geq 1\}}\frac{|u|^{p^{\ast}}}{|x|^{\gamma}}\,\mathrm{d}x$
$\leq\operatorname{\exp}(\alpha)\int_{\{x\in\mathbb{R}^{n}\colon~{}|x|<1\}}\frac{1}{|x|^{\gamma}}\,\mathrm{d}x+\operatorname{\exp}(\alpha)\int_{\{x\in\mathbb{R}^{n}\colon~{}|x|\geq 1\}}|u|^{p^{\ast}}\,\mathrm{d}x\leq C(\alpha,\gamma,n,p),$ (3.2)
where $C(\alpha,\gamma,n,p)>0$ is a constant depending only on $\alpha$, $\gamma$, $n$ and $p$. In order to evaluate $J_{1}$, we first denote $w(x):=u(x)-H(u)$ in $S$. It is easy to see that $w\in W^{2,\frac{n}{2}}_{\mathcal{N}}(S)$. As a result, we have the following estimates
$|u|^{\frac{n}{n-2}}$ $\leq(|w|+H(u))^{\frac{n}{n-2}}\leq|w|^{\frac{n}{n-2}}+\frac{n}{n-2}2^{\frac{n}{n-2}-1}\big{[}|w|^{\frac{n}{n-2}-1}H(u)+|H(u)|^{\frac{n}{n-2}}\big{]}$
@@
$\leq\bigg{[}1+\frac{2^{\frac{n}{n-2}}}{n-2}|H(u)|^{\frac{n}{2}}\bigg{]}|w|^{\frac{n}{n-2}}+D_{n},$
we thank to the well known Young’s inequality and the following elementary inequality of calculus
$(a+b)^{\wp}\leq a^{\wp}+\wp 2^{\wp-1}(a^{\wp-1}b+b^{\wp}),~{}\forall~{}\wp\geq 1;~{}a,b\geq 0.$
@@
Observe that $\xi\in W^{2,\frac{n}{2}}_{\mathcal{N}}(S)$ and thus $|u|^{\frac{n}{n-2}}\leq|\xi|^{\frac{n}{n-2}}+D_{n}$ in $S$. In addition, by direct calculation, we get
@@
$J_{1}$ $\leq\int_{S}\frac{\operatorname{\exp}(\alpha|u|^{\frac{n}{n-2}})}{|x|^{\gamma}}\,\mathrm{d}x\leq\int_{S}\frac{\operatorname{\exp}(\beta_{\gamma,n}(|\xi|^{\frac{n}{n-2}}+D_{n}))}{|x|^{\gamma}}\,\mathrm{d}x$
$\leq\operatorname{\exp}(\beta_{\gamma,n}D_{n})C(\gamma,n)|S|^{1-\frac{\gamma}{n}}\leq C(\gamma,n,p),$ (3.3)
where $C(\gamma,n,p)>0$ is a constant depending only on $\gamma$, $n$ and $p$. In virtue of (3) and (3), we infer that (1.13) holds.
Next, to check the validity of the sharpness of $\beta_{\gamma,n}$, inspired by Lu-Yang [39], we can define the following sequence
$\xi_{k}(x):=\begin{cases}\Big{(}\frac{\ln k}{\beta(n,2)}\Big{)}^{1-\frac{2}{n}}+(1-k^{\frac{2}{n}}|x|^{2})\frac{n\beta(n,2)^{\frac{2}{n}-1}}{2(\ln k)^{\frac{2}{n}}}\quad\text{if}~{}|x|\in[0,k^{-\frac{1}{n}}],\\ n\beta(n,2)^{\frac{2}{n}-1}(\ln k)^{-\frac{2}{n}}\ln\big{(}\frac{1}{|x|}\big{)}\quad\quad\quad\quad\quad\quad\text{if}~{}|x|\in[k^{-\frac{1}{n}},1],\\ \eta_{k}(x)\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\text{if}~{}|x|\in[1,+\infty),\end{cases}$ (3.4)
@@
Moreover, $\eta_{k}$ and $\Delta\eta_{k}$ are all of $O((\ln k)^{-\frac{2}{n}})$. By direct calculation, one can notice that
@@
Hence we have
$\|\Delta\xi_{k}\|_{p}^{p}$ =$=\int_{\{x\in\mathbb{R}^{n}\colon~{}1<|x|<2\}}|\Delta\eta_{k}|^{p}\,\mathrm{d}x+\omega_{n-1}\int_{0}^{k^{-\frac{1}{n}}}\big{(}n^{2}\beta(n,2)^{\frac{2}{n}-1}k^{\frac{2}{n}}(\ln k)^{-\frac{2}{n}}\big{)}^{p}r^{n-1}\,\mathrm{d}r$
@@
$=O((\ln k)^{-\frac{2p}{n}})+\frac{\omega_{n-1}}{nk}\bigg{(}n^{2}\beta(n,2)^{\frac{2}{n}-1}\bigg{(}\frac{k}{\ln k}\bigg{)}^{\frac{2}{n}}\bigg{)}^{p}+\frac{\omega_{n-1}}{n-2p}\big{(}n(n-2)\beta(n,2)^{\frac{2}{n}-1}(\ln k)^{-\frac{2}{n}}\big{)}^{p}\bigg{(}1-\frac{1}{k^{\frac{n-2p}{n}}}\bigg{)}$
$\leq O((\ln k)^{-\frac{2p}{n}})+\frac{\omega_{n-1}}{nk}\bigg{(}n^{2}\beta(n,2)^{\frac{2}{n}-1}\bigg{(}\frac{k}{\ln k}\bigg{)}^{\frac{2}{n}}\bigg{)}^{p}+\frac{\omega_{n-1}}{n-2p}\big{(}n(n-2)\beta(n,2)^{\frac{2}{n}-1}(\ln k)^{-\frac{2}{n}}\big{)}^{p}$
@@
$\|\Delta\xi_{k}\|_{\frac{n}{2}}^{\frac{n}{2}}$ =$=\int_{\{x\in\mathbb{R}^{n}\colon~{}1<|x|<2\}}|\Delta\eta_{k}|^{\frac{n}{2}}\,\mathrm{d}x+\omega_{n-1}\int_{0}^{k^{-\frac{1}{n}}}\big{(}n^{2}\beta(n,2)^{\frac{2}{n}-1}k^{\frac{2}{n}}(\ln k)^{-\frac{2}{n}}\big{)}^{\frac{n}{2}}r^{n-1}\,\mathrm{d}r$
@@
$=O((\ln k)^{-1})+\frac{\omega_{n-1}}{nk}\bigg{(}n^{2}\beta(n,2)^{\frac{2}{n}-1}\bigg{(}\frac{k}{\ln k}\bigg{)}^{\frac{2}{n}}\bigg{)}^{\frac{n}{2}}+\omega_{n-1}\big{(}n(n-2)\beta(n,2)^{\frac{2}{n}-1}(\ln k)^{-\frac{2}{n}}\big{)}^{\frac{n}{2}}\frac{\ln k}{n}$
@@
It follows that $\|\xi_{k}\|\to 1$ as $k\to\infty$. Let us assume $\alpha>\beta_{\gamma,n}$, then there exists $\delta>0$ such that $\alpha=(1+\delta)\beta_{\gamma,n}$. Consequently, one has
@@
$\geq\frac{\omega_{n-1}}{(n-\gamma)k^{\frac{n-\gamma}{n}}}~{}\Phi_{\alpha,j_{0}}\bigg{(}\frac{1}{\|\xi_{k}\|}\Big{(}\frac{\ln k}{\beta(n,2)}\Big{)}^{1-\frac{2}{n}}\bigg{)}$
$\sim\frac{\omega_{n-1}}{(n-\gamma)k^{\frac{n-\gamma}{n}}}~{}\operatorname{\exp}\bigg{(}\alpha\bigg{|}\frac{1}{\|\xi_{k}\|}\Big{(}\frac{\ln k}{\beta(n,2)}\Big{)}^{1-\frac{2}{n}}\bigg{|}^{\frac{n}{n-2}}\bigg{)}$
$=\frac{\omega_{n-1}}{(n-\gamma)}~{}\operatorname{\exp}\bigg{[}\bigg{(}1-\frac{\gamma}{n}\bigg{)}\ln k\bigg{(}\frac{1+\delta}{\|\xi_{k}\|^{\frac{n}{n-2}}}-1\bigg{)}\bigg{]}\to+\infty\quad\text{as}\quad k\to\infty,$
which completes the proof of Theorem 1.3. ∎
As a byproduct of the above theorem, we have the following corollary.
@@
Suppose that $1<p<\frac{n}{2}$, $0\leq\gamma<n$ and $n\geq 4$. Then for all $\alpha\geq 0$ and $u\in E$, there holds
@@
Since $C^{\infty}_{0}(\mathbb{R}^{n})$ in dense in $E$, for a given $\varepsilon>0$, one can assume $\psi\in C^{\infty}_{0}(\mathbb{R}^{n})$ such that $\|u-\psi\|<\varepsilon$. Define
@@
Notice that $|u(x)|\leq 1+\sup_{x\in\mathbb{R}^{n}}|\psi(x)|:=C$ for all $x\in S_{1}$, where $C>0$ is a suitable constant. It follows that $\Phi_{\alpha,j_{0}}(u)\lesssim|u|^{\frac{nj_{0}}{n-2}}$, where $\frac{nj_{0}}{n-2}\geq p^{\ast}$. Hence by using Lemma 2.5, we have
$\int_{S_{1}}\frac{\Phi_{\alpha,j_{0}}(u)}{|x|^{\gamma}}\,\mathrm{d}x\lesssim\int_{S_{1}}\frac{|u|^{\frac{nj_{0}}{n-2}}}{|x|^{\gamma}}\,\mathrm{d}x\lesssim\int_{\{x\in S_{1}:~{}|x|<1\}}\frac{1}{|x|^{\gamma}}\,\mathrm{d}x+\int_{\{x\in S_{1}:~{}|x|\geq 1\}}|u|^{\frac{nj_{0}}{n-2}}\,\mathrm{d}x<+\infty.$ (3.5)

### v3

@@ 2. Preliminary results
In this section, we shall discuss the basic properties of Schwarz symmetrization and some useful lemmas. In this regard, we refer to [7, 10]. First, we introduce the symmetrization of a Lebesgue measurable set $S$ in $\mathbb{R}^{n}$. Let $S^{\ast}$ be the open ball centered at the origin of $\mathbb{R}^{n}$ such that $|S^{\ast}|=|S|$. Further, we denote $|S|=0$ implies $S^{\ast}$ is the empty set, $|S|=+\infty$ implies $S^{\ast}=\mathbb{R}^{n}$ and $0<|S|<+\infty$ implies $S^{\ast}=B_{R}(0)$ such that $|S^{\ast}|=|S|=\sigma_{n}R^{n}$, where $\sigma_{n}$ is the volume of the unit ball in $\mathbb{R}^{n}$ and $R=R(S)$ is the volume radius of $S$. It is well-known that $S^{\ast}$ is called the symmetric decreasing rearrangement of $S$. Let $f\colon S\to\mathbb{R}$ be a measurable function vanishing at infinity, which means that
@@
Its distribution function $\mu_{f}\colon[0,+\infty)\to[0,+\infty]$ is given by $\mu_{f}(t)=|\{x\in S\colon~|f(x)|>t\}|$ for all $t\geq 0$. Note that $\mu_{f}$ is always a nonnegative, decreasing and right continuous function on $0,+\infty)$. Moreover, the unidimensional rearrangement of $f$ is a function $f^{\sharp}\colon[0,|S|]\to[0,+\infty]$, which is defined by
@@
where $m_{\mu_{f}}$ is the distribution function of $\mu_{f}$. Besides this, $f^{\sharp}$ is a nonnegative, decreasing and right continuous function on $0,+\infty)$. The maximal function of $f^{\sharp}$ is denoted by the symbol $f^{\sharp\sharp}$, and given by
(2.1) $f^{\sharp\sharp}(s):=\frac{1}{s}\int_{0}^{s}f^{\sharp}(t)\,\mathrm{d}t,~\forall~~s>0.$
In addition, $f^{\sharp\sharp}$ is nonnegative, decreasing and continuous on $(0,+\infty)$ and $f^{\sharp}\leq f^{\sharp\sharp}$. Now, we denote the function $f^{\ast}\colon S^{\ast}\to[0,+\infty]$ as the Schwarz symmetrization of $f$, alternatively known as the spherically symmetric decreasing rearrangement of $f$, and defined by
@@
In particular, $f^{\ast}$ is a radially symmetric, nonnegative and nonincreasing function. Moreover, if $\Psi\colon\mathbb{R}\to\mathbb{R}$ is nondecreasing, then we have
(2.2) $(\Psi(f))^{\ast}=\Psi(f^{\ast}).$
@@
In addition, we also have $\|f\|_{p}=\|f^{\ast}\|_{p}$ for all $p\in[1,+\infty)$ and the following Hardy-Littlewood inequality holds.
@@
[7, Corollary 2.16] Suppose that $f$ and $g$ are two nonnegative measurable functions on $\mathbb{R}^{n}$ with $\mu_{f}(t)<+\infty$ and $\mu_{g}(t)<+\infty$ for all $t>0$. Then there holds
(2.3) $\int_{\mathbb{R}^{n}}f(x)g(x)\,\mathrm{d}x\leq\int_{\mathbb{R}^{n}}f^{\ast}(x)g^{\ast}(x)\,\mathrm{d}x$
with the understanding that when the left side is infinite, so is the right side. Moreover, if $\int_{\mathbb{R}^{n}}f(x)g(x)\,\mathrm{d}x<+\infty$, then the equality in (2.3) holds if and only if $|\{(x,y)\in\mathbb{R}^{2n}\colon~f(x)<(f(y)~\text{and}~g(x)>g(y)\}|=0.$
@@
[44, Proposition 2.2] Let $u\in C^{\infty}_{0}(\mathbb{R}^{n})$ and $\Delta u=f\in C^{\infty}_{0}(\mathbb{R}^{n})$, then there exists a constant $D(n)$ depends upon only $n$ such that
@@
[11, Lemma A.VI] If $u\in L^{\wp}(\mathbb{R}^{n})$ for all $\wp\geq 1$ is a radial nonincreasing function, then there holds
@@
The following corollary follows from [52, Lemma 2.1 and Lemma 2.2].
@@
The function $s\mapsto\operatorname{\exp}(s)-\sum_{j=0}^{j_{0}-1}\frac{s^{j}}{j!}$ is continuous, increasing and convex in $0,+\infty)$, where $j_{0}$ is defined in (1.10). Moreover, for any $\alpha>0$, $\wp>1$ and $n\geq 4$, we have
@@
[25, Theorem 1.2 and Lemma 3.5] If $q\in[1,+\infty]$, $r\in[1,+\infty)$, $\in\mathbb{N}, j<kj<k$ such that $\frac{1}{p}=\frac{j}{kr}+\frac{k-j}{kq}$, then there exists a constant $C>0$ independent of $u$ such that
@@
3. Sharp singular Adams’ type inequality: Proof of Theorem 1.1
In this section, our aim is to prove Theorem 1.1. To establish Theorem 1.1, we shall use a rearrangement-free approach introduced by Lam-Lu [33]. Furthermore, as a consequence of Theorem 1.1 and more specifically inspired by Nguyen [44], we shall also demonstrate the $L^{1}$-integrability of the Young function (1.10) with singular weights on $\mathbb{R}^{n}$ in the desired function space.
@@
(3.1) $H(u)<1\quad\text{and}\quad|S|\leq 2^{\frac{4p^{\ast}}{n(n-2)}}C,$
where $C>0$ is a suitable constant depending only on $p$ and $n$, thanks to the chain of embeddings $E\hookrightarrow D^{2,p}(\mathbb{R}^{n})\hookrightarrow L^{p^{\ast}}(\mathbb{R}^{n})$. This shows that $S$ is a bounded domain. In addition, we can notice that
@@
(3.2) $J_{2}$ $\leq\int_{\{x\in\mathbb{R}^{n}\colon~u(x)<1\}}\frac{1}{|x|^{\gamma}}\sum_{j=j_{0}}^{\infty}\frac{\alpha^{j}}{j!}|u|^{\frac{nj}{n-2}}\,\mathrm{d}x\leq\int_{\{x\in\mathbb{R}^{n}\colon~u(x)<1\}}\frac{1}{|x|^{\gamma}}\sum_{j=j_{0}}^{\infty}\frac{\alpha^{j}}{j!}|u|^{p^{\ast}}\,\mathrm{d}x$
$\leq\operatorname{\exp}(\alpha)\Bigg[\int_{\{x\in\mathbb{R}^{n}\colon~u(x)<1,~|x|<1\}}\frac{|u|^{p^{\ast}}}{|x|^{\gamma}}\,\mathrm{d}x+\int_{\{x\in\mathbb{R}^{n}\colon~u(x)<1,~|x|\geq 1\}}\frac{|u|^{p^{\ast}}}{|x|^{\gamma}}\,\mathrm{d}x\Bigg]$
$\leq\operatorname{\exp}(\alpha)\Bigg[\int_{\{x\in\mathbb{R}^{n}\colon~|x|<1\}}\frac{1}{|x|^{\gamma}}\,\mathrm{d}x+\int_{\{x\in\mathbb{R}^{n}\colon~|x|\geq 1\}}|u|^{p^{\ast}}\,\mathrm{d}x\Bigg]$
$\leq C(\alpha,\gamma,p,n),$
where $C(\alpha,\gamma,p,n)>0$ is a constant depending only on $\alpha$, $\gamma$, $p$ and $n$. To evaluate $J_{1}$, we first denote w(x):=u(x)H(u)w(x):=u(x)-H(u) in $S$. It is easy to see that $w\in W^{2,\frac{n}{2}}_{\mathcal{N}}(S)$. As a result, we have the following estimates
$|u|^{\frac{n}{n-2}}$ $\leq(|w|+H(u))^{\frac{n}{n-2}}\leq|w|^{\frac{n}{n-2}}+\frac{n}{n-2}2^{\frac{n}{n-2}-1}\bigg[|w|^{\frac{n}{n-2}-1}H(u)+|H(u)|^{\frac{n}{n-2}}\bigg]$
@@
$\leq\bigg[1+\frac{2^{\frac{n}{n-2}}}{n-2}|H(u)|^{\frac{n}{2}}\bigg]|w|^{\frac{n}{n-2}}+D_{n}\quad\text{in}~~S,$
we thank to the well-known Young’s inequality and the following elementary inequality of calculus
$(a+b)^{\wp}\leq a^{\wp}+\wp 2^{\wp-1}(a^{\wp-1}b+b^{\wp}),~\forall~\wp\geq 1\quad\text{and}\quad a,b\geq 0.$
@@
Observe that $\xi\in W^{2,\frac{n}{2}}_{\mathcal{N}}(S)$ and thus, one has $|u|^{\frac{n}{n-2}}\leq|\xi|^{\frac{n}{n-2}}+D_{n}$ in $S$. In addition, by direct calculation, we get
@@
(3.3) $J_{1}$ $\leq\int_{S}\frac{\operatorname{\exp}(\alpha|u|^{\frac{n}{n-2}})}{|x|^{\gamma}}\,\mathrm{d}x\leq\int_{S}\frac{\operatorname{\exp}(\beta_{\gamma,n}(|\xi|^{\frac{n}{n-2}}+D_{n}))}{|x|^{\gamma}}\,\mathrm{d}x$
$\leq\operatorname{\exp}(\beta_{\gamma,n}D_{n})C(\gamma,n)|S|^{1-\frac{\gamma}{n}}\leq C(\gamma,p,n),$
where $C(\gamma,p,n)>0$ is a constant depending only on $\gamma$, $p$ and $n$. In virtue of (3.2) and (3.3), we infer that (1.11) holds. Next, to check the validity of the sharpness of $\beta_{\gamma,n}$, inspired by Lu-Yang [39], we can define the following sequence
(3.4) $\xi_{k}(x):=\begin{cases}\Big(\frac{\ln k}{\beta(n,2)}\Big)^{1-\frac{2}{n}}+(1-k^{\frac{2}{n}}|x|^{2})\frac{n\beta(n,2)^{\frac{2}{n}-1}}{2(\ln k)^{\frac{2}{n}}}\qquad\text{if}~|x|\in[0,k^{-\frac{1}{n}}],\\ n\beta(n,2)^{\frac{2}{n}-1}(\ln k)^{-\frac{2}{n}}\ln\big(\frac{1}{|x|}\big)\quad\quad\quad\quad\quad\quad\text{if}~|x|\in[k^{-\frac{1}{n}},1],\\ \eta_{k}(x)\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\text{if}~|x|\in[1,+\infty),\end{cases}$
@@
Moreover, $\eta_{k}$ and $\Delta\eta_{k}$ are all of $O((\ln k)^{-\frac{2}{n}})$. By direct calculations, one can notice that
@@
Hence, we have
$\|\Delta\xi_{k}\|_{p}^{p}$ $=\int_{\{x\in\mathbb{R}^{n}\colon~1<|x|<2\}}|\Delta\eta_{k}|^{p}\,\mathrm{d}x+\omega_{n-1}\int_{0}^{k^{-\frac{1}{n}}}\big(n^{2}\beta(n,2)^{\frac{2}{n}-1}k^{\frac{2}{n}}(\ln k)^{-\frac{2}{n}}\big)^{p}r^{n-1}\,\mathrm{d}r$
@@
$=O((\ln k)^{-\frac{2p}{n}})+\frac{\omega_{n-1}}{nk}\bigg(n^{2}\beta(n,2)^{\frac{2}{n}-1}\bigg(\frac{k}{\ln k}\bigg)^{\frac{2}{n}}\bigg)^{p}$
$\quad+\frac{\omega_{n-1}}{n-2p}\big(n(n-2)\beta(n,2)^{\frac{2}{n}-1}(\ln k)^{-\frac{2}{n}}\big)^{p}\bigg(1-\frac{1}{k^{\frac{n-2p}{n}}}\bigg)$
$\leq O((\ln k)^{-\frac{2p}{n}})+\frac{\omega_{n-1}}{nk}\bigg(n^{2}\beta(n,2)^{\frac{2}{n}-1}\bigg(\frac{k}{\ln k}\bigg)^{\frac{2}{n}}\bigg)^{p}$
$\quad+\frac{\omega_{n-1}}{n-2p}\big(n(n-2)\beta(n,2)^{\frac{2}{n}-1}(\ln k)^{-\frac{2}{n}}\big)^{p}$
@@
$\|\Delta\xi_{k}\|_{\frac{n}{2}}^{\frac{n}{2}}$ $=\int_{\{x\in\mathbb{R}^{n}\colon~1<|x|<2\}}|\Delta\eta_{k}|^{\frac{n}{2}}\,\mathrm{d}x+\omega_{n-1}\int_{0}^{k^{-\frac{1}{n}}}\big(n^{2}\beta(n,2)^{\frac{2}{n}-1}k^{\frac{2}{n}}(\ln k)^{-\frac{2}{n}}\big)^{\frac{n}{2}}r^{n-1}\,\mathrm{d}r$
@@
$=O((\ln k)^{-1})+\frac{\omega_{n-1}}{nk}\bigg(n^{2}\beta(n,2)^{\frac{2}{n}-1}\bigg(\frac{k}{\ln k}\bigg)^{\frac{2}{n}}\bigg)^{\frac{n}{2}}$
$\quad+\omega_{n-1}\big(n(n-2)\beta(n,2)^{\frac{2}{n}-1}(\ln k)^{-\frac{2}{n}}\big)^{\frac{n}{2}}\frac{\ln k}{n}$
@@
It follows that $\|\xi_{k}\|\to 1$ as $k\to\infty$. Let us assume $\alpha>\beta_{\gamma,n}$, then there exists $\delta>0$ such that $\alpha=(1+\delta)\beta_{\gamma,n}$. Moreover, one has
@@
$\geq\frac{\omega_{n-1}}{(n-\gamma)k^{\frac{n-\gamma}{n}}}~\Phi_{\alpha,j_{0}}\bigg(\frac{1}{\|\xi_{k}\|}\bigg(\frac{\ln k}{\beta(n,2)}\bigg)^{1-\frac{2}{n}}\bigg)$
$\sim\frac{\omega_{n-1}}{(n-\gamma)k^{\frac{n-\gamma}{n}}}~\operatorname{\exp}\Bigg(\alpha\Bigg|\frac{1}{\|\xi_{k}\|}\bigg(\frac{\ln k}{\beta(n,2)}\bigg)^{1-\frac{2}{n}}\Bigg|^{\frac{n}{n-2}}\Bigg)$
$=\frac{\omega_{n-1}}{(n-\gamma)}~\operatorname{\exp}\Bigg[\bigg(1-\frac{\gamma}{n}\bigg)\ln k\bigg(\frac{1+\delta}{\|\xi_{k}\|^{\frac{n}{n-2}}}-1\bigg)\Bigg]\to+\infty\quad\text{as}\quad k\to\infty.$
This completes the proof of Theorem 1.1. ∎
To end this section, we prove the following result, which is a byproduct of the previously discussed theorem.
@@
Let $1<p<\frac{n}{2}$, $0\leq\gamma<n$ and $n\geq 4$ be hold. Then for all $\alpha\geq 0$ and $u\in E$, there holds
@@
By employing the denseness of $C^{\infty}_{0}(\mathbb{R}^{n})$ in $E$, without loss of generality, for a given $\varepsilon>0$, one can assume $\psi\in C^{\infty}_{0}(\mathbb{R}^{n})$ such that $\|u-\psi\|<\varepsilon$. Moreover, we define two sets $S_{1}$ and $S_{2}$ as follows
@@
Notice that $|u(x)|\leq 1+\sup_{x\in\mathbb{R}^{n}}|\psi(x)|:=c$ for all $x\in S_{1}$, where $c>0$ is a suitable constant. By employing Lemma 2.5 and using the fact that $\frac{nj_{0}}{n-2}\geq p^{\ast}$, we have
(3.5) $\int_{S_{1}}\frac{\Phi_{\alpha,j_{0}}(u)}{|x|^{\gamma}}\,\mathrm{d}x$ $=\int_{S_{1}}\sum_{j=j_{0}}^{\infty}\frac{\big(\alpha c^{\frac{n}{n-2}}\big)^{j}}{j!}\bigg|\frac{u}{c}\bigg|^{\frac{nj}{n-2}}\,\frac{\mathrm{d}x}{|x|^{\gamma}}\leq\sum_{j=j_{0}}^{\infty}\frac{\big(\alpha c^{\frac{n}{n-2}}\big)^{j}}{j!}\int_{S_{1}}\bigg|\frac{u}{c}\bigg|^{\frac{nj_{0}}{n-2}}\,\frac{\mathrm{d}x}{|x|^{\gamma}}$
$\lesssim\int_{S_{1}}\frac{|u|^{\frac{nj_{0}}{n-2}}}{|x|^{\gamma}}\,\mathrm{d}x\lesssim\int_{\{x\in\mathbb{R}^{n}:~|x|<1\}}\frac{1}{|x|^{\gamma}}\,\mathrm{d}x+\int_{\{x\in\mathbb{R}^{n}:~|x|\geq 1\}}|u|^{\frac{nj_{0}}{n-2}}\,\mathrm{d}x<+\infty.$